OpenStudy (anonymous):

Use the distributive property to simplify the expression below 6x -10 / -2

OpenStudy (anonymous):

(6x - 10)/ -2 = -1/2 (6x - 10) Distribute -1/2 = -1/2(6x) - (-1/2) 10 You can finish from here?

OpenStudy (anonymous):

nop im really lost